OpenStudy (anonymous):

Calculate the discriminant to determine the number of real roots. y = x^2 + 3x + 9 How many real roots does the equation have?

hartnn (hartnn):

know how to calculate the discriminant ?

OpenStudy (anonymous):

I think so...

hartnn (hartnn):

Compare \(y= x^2 + 3x + 9\) with \(y=ax^2+bx+c\) find a,b,c

hartnn (hartnn):

then your discriminant is just \(\Large D = {b^2-4ac}\)

OpenStudy (anonymous):

Sorry, I don't get it.

hartnn (hartnn):

did you get how to find a,b,c values ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Shouldn't It be a=1 b=3 c=9

hartnn (hartnn):

 did you get now ?